OpenStudy (anonymous):

find the slope-intercept form of the line whose slope is 5 and that passes through the point (-5,11)

OpenStudy (anonymous):

Slope intercept form is expressed as y = mx + b. What you do in this situation is insert the variables you do know and solve for b, the y-intercept. 11 = 5(-5) + b 11 = -25 + b 11 + 25 = 36 = b So then, the slope-intercept form of the line is then y = 5x + 36

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Not a problem. Thank you for the medal!

OpenStudy (anonymous):

you're welcome. i got 5x+y=36.

OpenStudy (anonymous):

Don't forget to solve for y to put it in slope-intercept form (y = mx + b). Be careful with your multiplication/subtraction too—when the equation is solved for y, the "mx" or in this case should be positive, since we already know our slope, m, as 5. Once you solve for "b", the y-intercept, x and y sort of can be considered placeholders

OpenStudy (anonymous):

ok, thank you.